1/2

Lord George Gordon Byron ,  Lord George Gordon Byron ,
Sie wallt in Schönheit  She Walks in Beauty

tradukita de Adolf Seubert

Sie wallt in Schönheit wie die Nacht, Wenn wolkenlos die Sterne ragen; Was Hell und Dunkel lieblich macht, Wird dir ihr Aug’ und Antlitz sagen, Wo sich ein holder Licht entfacht, Als wenn die Himmel heiter tagen.

She walks in Beauty, like the Night Of cloudless climes and starry skies; And all that’s best of dark and bright Meet in her aspect and her eyes; Thus mellowed to the tender light Which Heaven to gaudy day denies –

Ein Schatten mehr, ein Strahl davon Würd’ jene hohen Reize schwächen, Die aus der Flechten Rabenton Und aus der Miene Zauber sprechen, Wo Lichtgedanken künden schon, Aus welchem reinen Haus sie brechen.

One shade the more – one ray the less Had half impaired the nameless grace Which waves in every raven tress Or softly lightens o’er her face Where thoughts serenely sweet express How pure – how dear their dwelling place!

Und auf der Wange, die so blüht, Und auf der Stirne steht geschrieben, Im Lächeln, das ins Herz uns glüht, Daß sie nur voll von Tugendtrieben, Daß mild und friedlich ihr Gemüth, Und Unschuld athme nur ihr Lieben.

And on that cheek, and o’er that brow, So soft – so calm – yet eloquent, The smiles that win, the tints that glow, But tell of days in goodness spent – A mind at peace with all below – A Heart – whose love is innocent!

…  …

2/2

Traduko de la Angla poemo “She Walks in Beauty” de Lord George Gordon Byron (∗1788-01-22 – †1824-04-18) en la Germanan de Adolf Seubert.

Arg-1109-2255 (2015-01-17 19:02:42)

Tiu ĉi poem-germanigo troviĝas en la retejo http: // gutenberg. spiegel. de/ buch/ gedichte-6666/ 2 . Pri la tradukinto vidu la retejon http: // de. wikisource. org/ wiki/ ADB: Seubert,_Adolf .

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1109-2238 (2015-01-17 17:34:42)

Tiu ĉi poemo troviĝas en http: // petercochran. files. wordpress. com/ 2009/ 03/ hebrew_melodies. pdf .

http://www.poezio.net/version?poem-id=1109&version-id=2255,2238